OpenStudy (anonymous):

give exact and approximate solutions to three decimal places x^2+14x+49=25

OpenStudy (helder_edwin):

u can use the cuadratic formula: given \(ax^2+bx+c=0\) then \[ \large x=\frac{-b\pm\sqrt{b^2-4ac}}{2a} \]

OpenStudy (helder_edwin):

do u know how to use this?

OpenStudy (anonymous):

I'm learning. The answer I come up with is -2,-12

OpenStudy (helder_edwin):

u got it right!

OpenStudy (anonymous):

aahhh...thx

OpenStudy (lgbasallote):

factor out both sides \[x^2 + 14x + 49 = 25\] \[(x+7)^2 = 25\] ge the square root of both sides \[x + 7 = \pm 5\] subtract 7 from both isdes \[x = \pm 5 - 7\] therefore x = -2; x = -12

OpenStudy (helder_edwin):

in this case what @lgbasallote did is quite more elegant. the formula is the "easiest" way, but can be "crude".

OpenStudy (lgbasallote):

i think fourth-dimensionally :DDD